### v2

@@ 2.2 Joint Liability Contracts with a Group Size of Two
Under a joint liability agreement, each farmer’s ex-ante expected profit $\pi_{i}$ can be written as
@@
We emphasize that, in this model, both farmers have a chance of achieving a successful harvest $\overline{Y}$ with a probability of $e^{2}$, which gives them a profit of $p\overline{Y}-w$. The probability of one farmer defaulting is given by $e(1-e)$, which means that farmer $i$ would receive his/her own surplus from a successful harvest minus the other farmer’s deficit, $p\overline{Y}+p\underline{Y}-2w$. If both farmers fail at the same time, they will pay all of their incomes to the lenders, which means that their profit will be zero. The detailed profit distributions are summarized conveniently in Table 1.
@@
One farmer succeeds, another farmer succeeds too. $e^{2}$ $p\overline{Y}-w$
One farmer succeeds, another farmer fails. $e(1-e)$ $p\overline{Y}+p\underline{Y}-2w$
One farmer fails, another farmer succeeds. $(1-e)e$ $p\underline{Y}-p\underline{Y}=0$
One farmer fails, another farmer fails. $(1-e)^{2}$ $p\underline{Y}-p\underline{Y}=0$
@@
Before setting the total objective (utility) function for the whole group in the model, we first list a number of exogenous constraints which we introduce to the corresponding optimization program.
@@
Assuming that the other borrower also succeeds in his/her project, this borrower will only be responsible for his/her own payment. In such a scenario the earnings plus the discounted cost of the next loan would at least be equal to the earnings of strategically defaulting.
[…]
@@
[…]
CONSTRAINT 4. The lender must be able to sustain the lending game, or at least break even, over the periods. Thus the expected repayment amount of each borrower has to be at least as large as $L(1+\epsilon)$, where $\epsilon$ is a risk-free rate. It is important to note that given that all constraints have been established under the assumption that the group will successfully repay the total loans to ensure their eligibility for subsequent loan rounds, we have not factored in the scenario where all borrowers within the group experience failure on the left hand side of Constraint 4.
[…]
Since $p\underline{Y}-w<0$, it is clear that Constraint 3 is tighter than Constraint 2. In the ensuing analysis, we will only consider Constraint 1, Constraint 3, and Constraint 4. Firstly, $\pi_{i}$ and $w$ have a strictly negative relationship as can be observed from Equation (2.1). So, to maximize $\pi_{i}$, $w$ has to be set at a minimum value. From Constraint 4, the minimum value of $w$ should be set at
@@
. Next we substitute $w=\frac{L(1+\varepsilon)}{1-(1-e)^{2}}$ into Constraint 1, and rewrite the equation as:
@@
To simplify the problem even further, we set $C(e)=\frac{1}{2}ce^{2}$ (where the cost parameter $c$ is used to scale the total cost) as the disutility function. The disutility function $C(e)=\frac{1}{2}ce^{2}$ is a commonly used function in economics to represent the cost or disutility associated with effort or work. See e.g. clark1996satisfaction and kahneman1997back. The function is quadratic, meaning that as an effort increases, the cost or disutility of that effort increases at an increasing rate. In the context of an individual agricultural farmer, it is reasonable to use a disutility function to acknowledge that while some effort may be necessary, there is a point at which additional effort becomes more and more unpleasant or burdensome. For instance, a farmer may be willing to put in an extra effort to cultivate crops or tend to a livestock, but as the required effort increases, the farmer may experience fatigue, stress, and reduced productivity, which can negatively impact their overall well–being. Furthermore, the quadratic form of the disutility function is also supported by empirical evidence on the relationship between effort and well–being. It is reasonable to assume that farmers will experience a phenomenon known as a diminishing marginal utility of effort, which implies that the additional unit of well-being gained by exerting more effort decreases as the level of effort increases.
@@ 2.3 Main Results
Now that we have listed the assumptions and set-up for the model, we proceed to state a number of key results expressed conveniently in the form of propositions. These propositions provide useful insights into the optimal individual–ESE scores, optimal group size, and loan ceiling constraints. PROPOSITION 1 (Optimal individual–ESE score). Finding an optimal ESE score for both the lenders and borrowers in an individual ESE–joint liability model is critical because it can help address the challenge of promoting sustainability in agricultural production while ensuring access to credit for small-scale farmers. By incorporating environmental and social factors into credit evaluation systems, it can incentivize farmers to adopt sustainable practices that can lead to positive environmental outcomes and long–term economic benefits. At the same time lenders can mitigate the risks associated with lending to farmers by incorporating sustainability factors into their credit evaluation process. To find the optimal individual–$ESE$ score (noted in formulas as $E$), we take a first-order partial derivative w.r.t $E$ of the ex-ante expected payoff function to obtain the following result
@@
After rearranging terms, we obtain an optimal individual-$ESE$ score (where $E$ is between 0 and 100) as
@@
PROPOSITION 2 (Collapsing two constraints on loan ceiling to a single constraint). The two loan ceiling constraints (Constraint 1 and Constraint 3) arise from analyzing the success or failure of the other co–loaners assuming their own success in the project, but most of the previous studies in this area have not explored the connection between these constraints. Notably, in rezaei2017optimal, they found that a feasible range of loan ceiling depends on the borrower’s discount rate $\delta$. Nevertheless, it is possible to reduce these two constraints to a single constraint in our model. Recall the two constraints on the loan ceiling in Equation (2.8) and Equation (2.11), labelled respectively as $L_{1}$ and $L_{2}$.
$L_{1}\leqslant\frac{p\overline{Y}+p\underline{Y}}{2(1+\varepsilon)}\left[1-(1-e)^{2}\right]$
$L_{2}\leqslant\frac{P\underline{Y}}{\frac{2(1+\varepsilon)}{1-(1-e)^{2}}-\delta}$
Next we show that $L_{1}$ is always larger than $L_{2}$, which means that we can only use $L_{2}$ as a final loan ceiling bound. It simplifies the problem and potentially makes it easier to analyze and solve the model. In addition, $L_{2}$ takes $\delta$ into consideration, which may include more information and additional key factors. This conclusion is illustrated in Figure 1.
@@
Proof: If $L_{1}\leqslant L_{2}$, then
@@
From the above results, if $L_{1}\leqslant L_{2}$, the discount factor $\delta$ is always greater than 1, which leads to a contradiction. This leads us to conclude that $L_{1}>L_{2}$. PROPOSITION 3 (A higher individual–ESE score leads to a higher loan ceiling). Next we examine the relationship between the individual–ESE score and the loan ceiling. Intuitively a borrower with a high-quality ESE score should be eligible for a higher loan amount, and therefore the loan ceiling should also be higher. Since $L_{1}$ is always greater than $L_{2}$, we will use $L_{2}$ to determine the loan ceiling, which defines the maximum value of $L_{2}$ to be the loan ceiling.
@@
Then we take a partial derivative of $L_{2}$ w.r.t. $E$ to obtain
@@
It is clear that the denominator is greater than 0. The numerator is also greater than 0, which can be illustrated by writing out the following result
@@
A borrower with a higher optimal ESE score, which means he/she has a better overall performance, will be perceived as less risky by the lender. As a result the borrower will be offered a higher loan ceiling, which is the maximum amount of loan that a lender is willing to provide to a borrower. In turn, if a borrower wants to obtain a larger loan amount, he/she may be motivated to improve his/her ESE performance.
@@ 2.4 Joint Liability Contracts with a Group Size of $n$
In this subsection we present an individual ESE–joint liability model which formalizes the assumption that the group of borrowers is collectively responsible for the entire group loan (i.e., $nw$). The expected utility of a borrower $i$ who adopts a repayment strategy during any period when they receive financing is specified as
@@
Constraints similar to those in the case of a group size of two are as follows.
  1. 1.
    There is at least one borrower who is successful at his/her project and capable of paying the entire group’s repayment amount.
@@
  2. 2.
    For each successful borrower, the payoff of strategically defaulting cannot be larger than the payoff of repaying and being refinanced.
@@
  3. 3.
    The lender must be able to sustain the lending game, or at least break even, over the periods. Thus the expected repayment amount of each borrower has to be at least as large as $L(1+\varepsilon)$, where $\varepsilon$ is the risk-free rate. It is again important to note that, given that all constraints have been established under the presumption that the group will successfully repay the total loans in order to ensure their eligibility for subsequent loan rounds, we have not factored in the scenario where all borrowers within the group experience failure on the left hand side of this constraint.
@@
    Similarly, we obtain $w\geqslant\frac{L(1+\varepsilon)}{1-(1-e)^{n}}$. Thus the value of $w$ should be set as $\frac{L(1+\varepsilon)}{1-(1-e)^{n}}$ due to the negative relationship between $w$ and $\pi_{i}$.
@@
PROPOSITION 4 (Optimal individual–ESE score and group size $n$ have a negative relationship). A higher ESE score indicates a lower default risk, which means that the borrower is more creditworthy and can obtain a loan with a smaller group size. On the other hand each member of the group is jointly responsible for the repayment of the loan, and if one member defaults, the other members will have to repay the outstanding amount. Thus a low-quality borrower may choose to loan with a larger group to reduce the probability of default and minimize the potential burden on any one individual. Next we verify this intuition based on the following derivations. We take a first-order partial derivative of the expected payoff w.r.t $E$ to obtain an optimal individual–ESE score for a borrower
$\frac{\partial\pi}{\partial E}=kp\overline{Y}+p\underline{Y}k\left[n(1-kE-b)^{n-1}-1\right]-k\frac{\partial c(e)}{\partial e}=0$ (2.25)
Since it is hard to find an explicitly analytical form of $n$ and $E$, we adopt an implicit function derivation below. Let $E(n)$ replace $E$ and take the partial derivative w.r.t. $n$ on both sides of $kp\overline{Y}+p\underline{Y}\left[n(1-kE-b)^{n-1}-1\right]-kc(kE+b)=0$. Then we obtain the following result
$p\underline{Y}\left\{(1-kE(n)-b)^{n-1}+n\cdot\frac{\partial\left[e^{(n-1)\log(1-kE(n)-b)}\right]}{\partial n}\right\}-C\left[k\frac{\partial E(n)}{\partial n}\right]=0$ (2.26)
@@
After reorganizing the formula, we can express the partial derivative of $E(n)$ w.r.t $n$ as
$\frac{\partial E(n)}{\partial n}=\frac{p\underline{Y}\left\{(1-e)^{n-1}[1+n\cdot\log(1-e)]\right\}}{p\underline{Y}\left[n\cdot(1-e)^{n-1}\cdot\frac{(n-1)k}{1-e}\right]+ck}$ (2.28)
It is clear that the denominator in Equation (2.28) is greater than 0. When $n\geqslant 1$ and $e\in(0,1)$, then $n\log(1-e)<-1$, and $(1-e)^{n-1}[1+n\cdot\log(1-e)]<0$. Thus, the numerator is less than 0. As a result $\frac{\partial E(n)}{\partial n}<0$, which proves the optimal individual–ESE score and group size $n$ have a negative relationship. In the light of the negative correlation between optimal individual–ESE scores and a group size, it is recommended that insurance companies prioritize lenders with higher ESE scores when the group size is smaller to mitigate potential risks. Borrowers with high scores are typically perceived to exhibit positive ESE performance, which leads to higher expectations for sustained success and repayment probability over the long term. Despite the relatively small group size, lenders may still have heightened confidence in the ability of high–scoring borrowers to demonstrate honesty and success, and to exhibit willingness to repay their loans. As a result lenders may consider reducing group size requirements for high-scoring borrower groups. Conversely, as the group size increases, the requirement for lenders’ ESE scores may potentially decrease, as a larger group may possess a greater risk mitigation capacity and an increased likelihood of successfully repaying the group loan. Consequently lenders may opt to reduce the score requirements in such cases. PROPOSITION5 (Optimal individual–ESE score converges to a constant level for a large enough group size). Equation (2.28) tells us that $\lim_{n\rightarrow\infty}\frac{\partial E(n)}{\partial n}=0$, which means that the optimal individual–ESE score converges to a constant level when the group size reaches a certain large enough number. To find the limit of the optimal score value, we recall the first-order condition of the ex-ante expected payoff, which is
@@
$\frac{\partial C(e)}{\partial e}=ce=c(kE+b).$
@@
$E\rightarrow\left[\frac{p(\overline{Y}-\underline{Y})}{c}-b\right]/k$ (2.30)
To illustrate this proposition, we provide a numerical example, which is shown in Figure 2. We set the unit price $p=1$, the successful probability $e=k\times E+b=1/100\times E+0$, high yield amount $\overline{Y}=1000$, low yield amount $\underline{Y}=500$, and perform a simulation under the different group size $n\in[1,100]$.
@@
In situations where the group size is small, insurance companies may seek to attract borrowers with high ESE scores in order to mitigate risk. However, as the group size expands, the marginal benefit of an individual effort on group outcomes diminishes, resulting in a reduction of an individual farmer’s optimal effort level. This can create incentives for farmers to contribute less effort, leading to a ‘free-rider’ problem and decreased profits for the group. Therefore the threshold value of the optimal $ESE$ score can serve as a reference point for both the lenders and borrowers, enabling them to avoid an sub-optimal resource allocation and a free-rider problem.

### v3

@@ 3.2 Joint Liability Contracts with a Group Size of Two
Under a joint liability agreement with a group size of two (‘Farmer A’ and ‘Farmer B’), from the view of ‘Farmer A’, his/her ex-ante expected profit $\pi_{i}$ can be written as
@@
We emphasize that, in this model, both farmers have a chance of achieving a successful harvest $\overline{Y}$ with a probability of $e^{2}$, which gives them a profit of $p\overline{Y}-w$. The probability of one farmer defaulting is given by e(1e)e(1-e), which means that farmer $i$ would receive his/her own surplus from a successful harvest minus the other farmer’s deficit, $p\overline{Y}+p\underline{Y}-2w$. If both farmers fail in their respective projects at the same time, they will pay all of their incomes to the lenders, which means that their profit will be zero. The detailed profit distributions are summarized conveniently in Table 3. The probabilities and profit outcomes are calculated entirely from the perspective of Farmer A, reflecting the cooperative repayment structure under the assumption that high-yield borrowers (successful farmers) repay their loans in full, while low-yield borrowers (unsuccessful farmers) repay only the income they earn ($p\underline{Y}$). When Farmer A succeeds ($p\overline{Y}$) and Farmer B fails, Farmer A must not only repay their own loan obligation $(w)$ but also cover the shortfall left by Farmer B, who repays only $p\underline{Y}$. This results in Farmer A’s profit being $p\bar{Y}-w-(w-p\underline{Y})=p\overline{Y}+p\underline{Y}$- $2w$. Conversely, when Farmer A fails and Farmer B succeeds, Farmer A’s repayment is limited to their income $(p\underline{Y})$, with Farmer B’s success ensuring the cooperative repayment obligation is met, leading to a profit for Farmer A of $p\underline{Y}-p\underline{Y}=0$. The asymmetry in profits arises from this risk-sharing mechanism, where successful farmers are expected to shoulder additional responsibility to maintain the system’s financial integrity, while unsuccessful farmers contribute as much as their income allows them to do so, shielding them from excessive financial burdens.
@@
‘Farmer A’ succeeds, ‘Farmer B’ succeeds too. $e^{2}$ $p\overline{Y}-w$
‘Farmer A’ succeeds, ‘Farmer B’ fails. e(1e)e(1-e) $p\overline{Y}+p\underline{Y}-2w$
‘Farmer A’ fails, ‘Farmer B’ succeeds. (1e)e(1-e)e $p\underline{Y}-p\underline{Y}=0$
‘Farmer A’ fails, ‘Farmer B’ fails. $(1-e)^{2}$ $p\underline{Y}-p\underline{Y}=0$
@@
Before setting the total objective (utility) function for the whole group in the model, we first list a number of exogenous constraints which we introduce to the optimization program.
The constraints used here are motivated by the commonly accepted assumptions in the literature on joint liability models. For example, Rezaei et al. (2017) introduced similar constraints in their analysis of the optimal group size in the joint liability contracts to ensure the feasibility and sustainability of the contract. These constraints are based on expected returns and the repayment strategies of group members. While our study adopts a similar structure for these constraints, the detailed settings of the expected profit and the underlying model assumptions differ from their study. Moreover, our model explicitly incorporates sustainability features, addressing dynamic adjustments and long-term stability of joint liability contracts under varying economic conditions and group compositions.
@@
Assuming that the other borrower also succeeds in his/her project, this borrower will only be responsible for his/her own payment. In this situation, the earnings plus the discounted cost of the next loan would at least be equal to the earnings of strategically defaulting.
$\text{Constraint 2: }p\overline{Y}\leqslant p\overline{Y}-w+\delta L$ (3.3)
@@
$\text{Constraint 3: }p\overline{Y}\leqslant p\overline{Y}+p\underline{Y}-2w+\delta L$ (3.4)
CONSTRAINT 4. The lender must be able to sustain the lending game, or at least break even, over the periods. Thus the expected repayment amount of each borrower has to be at least as large as $L(1+\epsilon)$, where $\epsilon$ is a risk-free rate. It is important to note that given that all constraints have been established under the assumption that the group will successfully repay the total loans to ensure their eligibility for subsequent loan rounds, we have not factored in the case, in which all borrowers within the group experience failure in their projects, on the left-hand side of Constraint 4.
$\text{Constraint 4: }w\left[1-(1-e)^{2}\right]\geqslant L(1+\varepsilon)$ (3.5)
Since $p\underline{Y}-w<0$, it is clear that Constraint 3 is tighter than Constraint 2. In the ensuing analysis, we will only consider Constraint 1, Constraint 3, and Constraint 4. Firstly $\pi_{i}$ and $w$ have a strictly negative relationship as can be observed from Equation (3.1). So, to maximize $\pi_{i}$, $w$ has to be set at a minimum value. From Constraint 4, the minimum value of $w$ should be set at
@@
Next we substitute $w=\frac{L(1+\varepsilon)}{1-(1-e)^{2}}$ into Constraint 1, and rewrite the equation as
@@
To simplify the problem even further, we set $C(e)=\frac{1}{2}ce^{2}$ (where the cost parameter $c$ is used to scale the total cost) as the disutility function. The disutility function $C(e)=\frac{1}{2}ce^{2}$ is a commonly used function in Economics to represent the cost or disutility associated with effort or work. See e.g. Clark and Oswald (1996) and Kahneman et al. (1997). The function is chosen to be quadratic so that it can accommodate the following relationship: as an effort increases, the cost or disutility of that effort increases at an increasing rate. In the context of an individual agricultural farmer, it is reasonable to use a disutility function to acknowledge that while some effort may be necessary, there is a point at which additional effort becomes more and more unpleasant or burdensome. For instance, a farmer may be willing to put in extra effort to cultivate crops or tend to livestock, but as the required effort increases, the farmer may experience fatigue, stress, and reduced productivity, which can negatively impact their overall well–being. Furthermore, the quadratic form of the disutility function is also supported by empirical evidence on the relationship between effort and well-being. It is reasonable to assume that farmers will experience a phenomenon known as a diminishing marginal utility of effort, which implies that the additional unit of well-being gained by exerting more effort decreases as the level of effort increases.
@@ 3.3 Main Results
In this subsection, we state a number of key results expressed conveniently in the form of propositions. These propositions provide useful insights into the optimal individual–ESE scores, optimal group size, and loan ceiling constraints.
PROPOSITION 1 (Optimal individual–ESE score). Finding an optimal ESE score for both the lenders and borrowers in an individual ESE–joint liability model is critical because it can help address the challenge of promoting sustainability in agricultural production while ensuring access to credit for small-scale farmers. By incorporating environmental and social factors into credit evaluation systems, it can incentivize farmers to adopt sustainable practices that can lead to positive environmental outcomes and long-term economic benefits. At the same time, lenders can mitigate the risks associated with lending to farmers by incorporating sustainability factors into their credit evaluation process.
To find the optimal value of individual–ESE score, $E$, we take a first-order partial derivative with respect to $E$ of the ex-ante expected payoff function to obtain the following result
@@
After rearranging terms, we obtain an optimal individual-ESE score (where $0\leqslant E\leqslant 100$) as
@@
PROPOSITION 2 (Collapsing two constraints on loan ceiling to a single constraint). The two loan ceiling constraints, Constraint 1 and Constraint 3, originate from analyzing different worst-case scenarios to ensure the affordability and sustainability of the group’s repayment structure. However, most of the previous studies in this area have not explored the connection between these two constraints. Notably, in Rezaei et al. (2017), they found that a feasible range of loan ceilings depends on the borrower’s discount rate $\delta$. Nevertheless, it is possible to reduce these two constraints to a single constraint in our model.
Recall the two constraints on the loan ceiling in Equation (3.8) and Equation (3.11), labelled respectively as $L_{1}$ and $L_{2}$. Next, we show that $L_{1}$ is always larger than $L_{2}$, which means that we can only use $L_{2}$ as a final loan ceiling bound. It simplifies the problem and potentially makes it easier to solve the model. In addition, $L_{2}$ takes $\delta$ into consideration, which may include more information and additional key factors. This result is illustrated in Figure 1.
@@
Proof:
If $L_{1}\leqslant L_{2}$, then
@@
From the above results, if $L_{1}\leqslant L_{2}$, the discount factor $\delta$ is always greater than 1, which leads to a contradiction. This leads us to conclude that $L_{1}>L_{2}$.
PROPOSITION 3 (A higher individual–ESE score leads to a higher loan ceiling). Next, we examine the relationship between the individual–ESE score and the loan ceiling. Intuitively a borrower with a high-quality ESE score should be eligible for a higher loan amount, and therefore the loan ceiling should also be higher.
Since $L_{1}$ is always greater than $L_{2}$, we will use $L_{2}$ to determine the loan ceiling, which defines the maximum value of $L_{2}$ to be the loan ceiling.
@@
Then we take a partial derivative of $L_{2}$ with respect to $E$ to obtain
@@
It is clear that the denominator is greater than 0. The numerator is also greater than 0, which can be illustrated more clearly by writing out the following expression
@@
A borrower with a higher optimal ESE score, which means that he/she has a better overall performance, will be perceived as less risky by the lender. As a result the borrower will be offered a higher loan ceiling, which is the maximum amount of loan that a lender is willing to provide to a borrower. In turn, if a borrower wants to obtain a larger loan amount, he/she may be motivated to improve his/her ESE performance.
@@ 3.4 Joint Liability Contracts with a Group Size of $n$
In this subsection, we present an individual ESE–joint liability model which formalizes the assumption that the group of borrowers is collectively responsible for the entire group loan (i.e., $nw$). The expected utility of a borrower $i$ who adopts a repayment strategy during any period when they receive financing is specified as
@@
Constraints similar to those in the case of a group size of two are as follows (where g stands for ‘general’).
CONSTRAINT 1g. There is at least one borrower who is successful at his/her project and capable of paying the entire group’s repayment amount.
@@
CONSTRAINT 2g. Since Constraint 3 is more restrictive than Constraint 2g when the group size is two, we will omit Constraint 2g from further consideration for a group size of $n$.
CONSTRAINT 3g. For each successful borrower, the payoff of strategically defaulting cannot be larger than the payoff of repaying and being refinanced.
@@
CONSTRAINT 4g. The lender must be able to sustain the lending game, or at least break even, over the periods. Thus the expected repayment amount of each borrower has to be at least as large as $L(1+\varepsilon)$, where $\varepsilon$ is the risk-free rate. As in the two-borrowers case, it is important to note that, given that all of the constraints have been established under the presumption that the group will successfully repay the total loans to ensure their eligibility for subsequent loan rounds, we have not factored in the case, where all borrowers within the group experience failure, on the left-hand side of this constraint.
@@
Similarly we obtain $w\geqslant\frac{L(1+\varepsilon)}{1-(1-e)^{n}}$. Thus the value of $w$ should be set as $\frac{L(1+\varepsilon)}{1-(1-e)^{n}}$ due to the negative relationship between $w$ and $\pi_{i}$.
@@
PROPOSITION 4 (Optimal individual–ESE score and group size $n$ have a negative relationship). A higher ESE score indicates a lower default risk, which means that the borrower is more creditworthy and can obtain a loan with a smaller group size. On the other hand, each member of the group is jointly responsible for the repayment of the loan, and if one member defaults, the other members will have to repay the outstanding amount. Thus, a low-quality borrower may choose to loan with a larger group to reduce the probability of default and minimize the potential burden on any one individual. Next, we verify this intuition based on the following derivations.
We take a first-order partial derivative of the expected payoff with respect to $E$ to obtain an optimal individual–ESE score for a borrower as below
$\frac{\partial\pi}{\partial E}=kp\overline{Y}+p\underline{Y}k\left[n(1-kE-b)^{n-1}-1\right]-k\frac{\partial C(e)}{\partial e}=0,$ (3.25)
where the cost is $C(e)=\frac{1}{2}ce^{2}$.
Since it is hard to find an explicitly analytical form of $n$ and $E$, we adopt an implicit function derivation below. Let $E(n)$ replace $E$ and take the partial derivative with respect to $n$ on both sides of $kp\overline{Y}+p\underline{Y}\left[n(1-kE-b)^{n-1}-1\right]-kc(kE+b)=0$. Then we obtain the following result
$p\underline{Y}\left\{(1-kE(n)-b)^{n-1}+n\cdot\frac{\partial\left[e^{(n-1)\log(1-kE(n)-b)}\right]}{\partial n}\right\}-c\left[e\frac{\partial E(n)}{\partial n}\right]=0$ (3.26)
@@
After reorganizing the formula, we can express the partial derivative of $E(n)$ with respect to $n$ as
$\frac{\partial E(n)}{\partial n}=\frac{p\underline{Y}\left\{(1-e)^{n-1}[1+n\cdot\log(1-e)]\right\}}{p\underline{Y}\left[n\cdot(1-e)^{n-1}\cdot\frac{(n-1)k}{1-e}\right]+ce}$ (3.28)
It is clear that the denominator in Equation (3.28) is greater than 0. When $n\geqslant 1$ and $e\in(0,1)$, then $n\log(1-e)<-1$, and $(1-e)^{n-1}[1+n\cdot\log(1-e)]<0$. Thus, the numerator is less than 0. As a result $\frac{\partial E(n)}{\partial n}<0$, which shows that the optimal individual–ESE score and group size $n$ have a negative relationship.
In the light of the negative correlation between the optimal individual–ESE scores and the group size, it is recommended that insurance companies prioritize lenders with higher ESE scores when the group size is smaller to mitigate potential risks. Borrowers with high scores are typically perceived to exhibit positive ESE performance, which leads to higher expectations for sustained success and repayment probability over the long term. Despite the relatively small group size, lenders may still have heightened confidence in the ability of high-scoring borrowers to demonstrate honesty and success and to exhibit willingness to repay their loans. As a result, lenders may consider reducing group size requirements for high-scoring borrower groups. Conversely, as the group size increases, the requirement for lenders’ ESE scores may potentially decrease, as a larger group may possess a greater risk mitigation capacity and an increased likelihood of successfully repaying the group loan. Consequently, lenders may opt to reduce the score requirements in such cases.
PROPOSITION 5 (Optimal individual–ESE score converges to a constant level for a large enough group size). Equation (3.28) tells us that $\lim_{n\rightarrow\infty}\frac{\partial E(n)}{\partial n}=0$, which means that the optimal individual–ESE score converges to a constant level when the group size reaches a certain large enough number. To find the limit of the optimal score value, we recall the first–order condition of the ex-ante expected payoff, which is
@@
$\frac{\partial C(e)}{\partial e}=ce=c(kE+b).$ (3.30)
@@
$E\rightarrow\left[\frac{p(\overline{Y}-\underline{Y})}{c}-b\right]/k.$ (3.31)
To illustrate this proposition, we provide a numerical example, which is shown in Figure 2. We set the unit price $p=1$, the successful probability $e=k\times E+b=1/100\times E+0$, high yield amount $\overline{Y}=1000$, low yield amount $\underline{Y}=500$, and perform a simulation under a different group size $n\in[1,100]$.
@@
In situations where the group size is small, insurance companies may seek to attract borrowers with high ESE scores in order to mitigate risk. However, as the group size expands, the marginal benefit of an individual effort on group outcomes diminishes, resulting in a reduction of an individual farmer’s optimal effort level. This can create incentives for farmers to contribute less effort, leading to a ‘free-rider’ problem and decreased profits for the group. Therefore the threshold value of the optimal ESE score can serve as a reference point for both the lenders and borrowers, enabling them to avoid a sub-optimal resource allocation and a ‘free-rider’ problem.
In the analysis above, $n\rightarrow\infty$ serves as a theoretical construct to explore the asymptotic behavior of the optimal ESE score, $E$. However, in practical applications, $n$, representing the group size, is bounded by real-world constraints. Factors such as coordination challenges, logistical limitations, and the effectiveness of group dynamics impose natural upper limits on the group size. From the results of the simulation plot, it is evident that the optimal ESE score converges once $n$ exceeds a certain threshold. Beyond this point, increasing the group size further does not contribute meaningfully to improving the ESE score and may instead exacerbate issues such as the ‘free-rider’ problem. Consequently, the results derived for large $n$ should be interpreted as indicative of the system’s behavior in scenarios where $n$ approaches the upper bounds of realistic group sizes. This bounded nature of $n$ ensures the applicability and relevance of the model to real-world contexts.